Match the H-NMR spectra to the compound. NMR 1 NMR 2 OH 8 7 6 5 4 PPM 65 PPM Premise Response Drag and drop to ma
Answer #1

o-chz CH₂o-H · Aryl Ô came at 7-8 ppm. • -o-chz = 3.5-4.0 ppm 50 we will differentiatten -O-CH, = 4.0-4.5ppm our NMR peakby-0